OpenStudy (anonymous):

60 per cent of what number is equal to 45 per cent of 80?

OpenStudy (campbell_st):

well if you look at it \[\frac{60}{100} \times a = \frac{45}{100} \times 80\] which is the same as \[0.6a = 36\] divide both sides of the equation by 0.6 to find a